# BlendScape: Enabling Unified and Personalized Video-Conferencing Environments through Generative AI

### v1

@@ Abstract.
Today’s video-conferencing tools support a rich range of professional and social activities, but their generic, grid-based environments cannot be easily adapted to meet the varying needs of distributed collaborators. To enable end-user customization, we developed BlendScape, a system for meeting participants to compose video-conferencing environments tailored to their collaboration context by leveraging AI image generation techniques. BlendScape supports flexible representations of task spaces by blending users’ physical or virtual backgrounds into unified environments and implements multimodal interaction techniques to steer the generation. Through an evaluation with 15 end-users, we investigated their customization preferences for work and social scenarios. Participants could rapidly express their design intentions with BlendScape and envisioned using the system to structure collaboration in future meetings, but experienced challenges with preventing distracting elements. We implement scenarios to demonstrate BlendScape ’s expressiveness in supporting distributed collaboration techniques from prior work and propose composition techniques to improve the quality of environments.
@@ 1. Introduction
Distributed collaboration tools are becoming essential to the way we conduct daily activities. Since the pandemic, there has been a rise in popularity of using video-conferencing tools for a rich set of collaboration scenarios, ranging from professional tasks (e.g., remote work and distance learning (Benabdallah et al., 2021; Lee et al., 2022)) to social gatherings (e.g., connecting with family members (Fuchsberger et al., 2021), practicing music (Cai et al., 2021), playing tabletop games (Yuan et al., 2021)). However, today’s video-conferencing tools do not reflect this wide variety of activities due to how these tools compose meeting environments (i.e., the “stage” or background rendered around users’ videos). Users are typically placed in separate regions of a video grid within generic rooms that lack a connection to the meeting context, which can lead to meeting fatigue (Fauville et al., 2021), reduce user engagement (Choi and Diehm, 2021), and disrupt interpersonal cues necessary for mediating conversations (Tang et al., 2023; Hyrkas et al., 2023; Hu et al., 2023a).
In recent video-mediated communication and HCI systems research, we observe three approaches to reimagining the composition of video-conferencing environments to better support distributed collaboration: (1) Establishing a shared context, e.g., through rendering users’ videos in a unified environment (O’Hara et al., 2011; Morikawa and Maesako, 1998) and setting the mood or theme of a meeting (Hunter et al., 2014; Follmer et al., 2010); (2) Enabling spatial metaphors for communication (e.g., by incorporating proxemic interactions (Hu et al., 2023a) and spatial sound (Hyrkas et al., 2023; Tang et al., 2023)) to help users mediate turn-taking); (3) Using the environment to record a meeting history, to provide a basis for later collaboration (e.g., through immersive authoring capabilities (Xia et al., 2018)). However, it is still a challenge for end-users to adapt meeting environments to meet their distributed collaboration needs during live meetings (Grønbæk et al., 2023a). Customization tools in commercial systems (e.g., Microsoft Teams Together Mode111Microsoft Teams TogetherMode: https://www.microsoft.com/en-us/microsoft-teams/teams-together-mode, Ohyay222Ohyay: https://ohyay.co/) require significant time and effort, which may lead users to default to generic, pre-defined options.
@@
In this paper, we present BlendScape, a rendering and composition system that enables video-conference participants to easily create environments tailored to the meeting context, with two key innovations: (1) We ground the generation of meeting environments in real spaces that are meaningful to users by blending their physical or virtual backgrounds into a unified environment. This can serve as a mechanism for personalization (Venolia et al., 2018) or to incorporate physical objects to collaborate around (Hu et al., 2023b; Junuzovic et al., 2012). (2) Capitalizing on recent advances in generative AI, we leverage image generation models to enable expressive and rapid techniques for composing environment designs. While such techniques are the subject of ongoing research, several dominant modes of use have emerged:
text-to-image: An image is generated from a given text prompt.
@@
image-to-image: An image is generated from a text prompt and an input image, retaining features of the input image while introducing new elements or styles consistent with the prompt.
@@
inpainting: Similar to image-to-image, but where an image mask is used to determine which parts of the input image should be unchanged. The rest of the image is generated in a way that it is consistent with the fixed parts of the input image (i.e., blended).
BlendScape uses (1) an inpainting technique that creates a new environment by merging users’ webcam backgrounds into a unified environment; (2) an image-to-image technique that transforms an existing image of an environment (i.e., an image prior) to reflect the meeting purpose and composites users’ videos within it. We enable meeting participants to steer the generation of environments through multimodal interaction techniques, including text-based prompts, direct manipulation, and choice of image priors. To lower the barrier for end-users to generate good quality meeting spaces, we developed composition techniques in BlendScape to guide the AI models to produce contextually relevant environments and integrate users’ videos within the scene in an immersive manner.
We evaluated BlendScape in two steps. First, to demonstrate the expressiveness of BlendScape’s composition techniques (Ledo et al., 2018), we implemented and analyzed three scenarios using BlendScape. These scenarios, guided by a review of prior work, incorporate a range of seven techniques for supporting distributed collaboration (e.g., conveying physical and digital task spaces, enabling ad-hoc layouts of users that reflect the meeting context, or recording a history of collaboration within a meeting space). Our analysis focused on the extent to which BlendScape can be used to implement these techniques in familiar or new ways.
@@
Second, we conducted a user study with 15 end-users to investigate their preferences for customizing environments for various meeting scenarios and to what extent BlendScape enables them to achieve their design intentions. Study participants compared and critiqued pre-generated environments for two social gatherings, then used BlendScape to design a series of meeting spaces for a remote work scenario. The participants found value in BlendScape’s composition techniques for enabling them to rapidly prototype environment designs for a wide range of meeting activities. However, they experienced challenges with mitigating distracting and unrealistic elements to achieve the ideal design. We propose extensions and improvements to BlendScape’s implementation to address these limitations in future work.
Our work contributes: (1) the BlendScape system, which enables end-user customization of video-conferencing environments through generative AI-driven composition techniques; (2) three scenarios that demonstrate how BlendScape can be used to prototype a range of seven techniques for supporting distributed collaboration from prior work; (3) an evaluation with 15 frequent users of video-conferencing tools, which surfaced benefits and limitations of BlendScape for creating customized meeting spaces.
@@ 2. Related Work
Our work builds on prior research on (1) composing video-conferencing environments to support distributed collaboration; (2) generative AI techniques for constructing 2D and 3D environments.
2.1. Video-Conferencing Environments
To situate our work in the video-mediated communication literature, we first classified techniques for composing meeting environments, based on a review of commercial video-conferencing tools and HCI systems research. Here, we define the environment as the “stage” or background rendered around users’ videos, excluding other functional elements of video-conferencing tools such as the chat or host controls. As shown in Fig. 2, we further categorized these composition techniques according to three overarching benefits they can provide for distributed collaboration, as demonstrated through empirical studies in prior work: (1) establishing a shared context, (2) enabling spatial metaphors for communication, (3) serving as a record or artifact of collaboration.
Classification of Environment Composition Techniques. This figure shows a table with three columns. The first column has header ”Establishing a Shared Context” and children ”Rendering users’ videos within a unified environment”, ”Conveying the physical or digital task space”, ”Setting the theme or mood of a meeting. The second column as header ”Enabling Spatial Metaphors for Communication” and children ”Enabling proxemic metaphors”, ”Enabling scaling metaphors”, ”Rendering spatialized audio”, ”Establishing consistent user layouts”, ”Rendering unique viewing perspectives.” The third column has header ”Serving as a Record or Artifact of Collaboration” and children ”Recording a history of collaborative activities”, ”Serving as an artifact produced through collaboration.”
We focus our review on screen-based meeting systems with 2D or 2.5D designs; composition techniques for other modalities (e.g., tabletop (Zagermann et al., 2016; Tuddenham and Robinson, 2009) or mixed reality interfaces (Pejsa et al., 2016; Herskovitz et al., 2022; Grønbæk et al., 2023b)) would likely introduce new dimensions to our categorization.
@@ 2.1. Video-Conferencing Environments
A primary role of virtual meeting environments is creating a shared frame of reference for distributed collaborators. One approach popularized by commercial tools (e.g., Teams Together Mode, Ohyay, Zoom Immersive View333Zoom Immersive View: https://blog.zoom.us/introducing-zoom-immersive-view/) as well as research systems (e.g., Waazam (Hunter et al., 2014), HyperMirror (Morikawa and Maesako, 1998), BISi (O’Hara et al., 2011)) is rendering users’ videos within a single, unified environment to simulate a co-located meeting. Prior studies found this approach promising for enhancing user engagement and increasing co-presence, as compared to traditional grid-based video layouts (Tang et al., 2023; Hunter et al., 2014). Teams and Zoom provide lightweight tools for uploading custom environment images and manually placing participants, whereas Ohyay offers an extensive customization suite for meeting organizers to design multi-room environments and script interactive behaviors. Yet, the process of tailoring environments to align with the meeting context can be highly manual and time-consuming with these tools, which may cause users to default to pre-defined templates.
Prior meeting tools have explored a variety of environment design techniques for conveying the task space, i.e., the physical or digital space where work takes place and artifacts are collaboratively produced (Buxton, 2009). Commercial tools like Teams and Zoom enable screen-sharing capabilities to display digital task spaces (e.g., slideshows, shared documents), which are rendered alongside users’ videos in a grid-based layout. MirrorBlender (Grønbæk et al., 2021) enables blending the person and task space through “mirrors” (i.e., translucent representations of users’ video feeds and shared screens) that can be layered on top of each other to enable more natural interaction techniques for hybrid collaboration (e.g., using hand gestures to interact with shared content). Conveying physical task spaces (e.g., for learning hands-on skills (Kumaravel et al., 2019) or practicing music (Cai et al., 2021)) is an ongoing challenge with commercial video-conferencing tools, often requiring users to implement custom hardware setups to capture their physical surroundings (Junuzovic et al., 2012; Labrie et al., 2022). To mitigate this, ThingShare (Hu et al., 2023b) and VideoPlay (Follmer et al., 2010) enable users to scan physical objects and manipulate them as digital representations during video calls.
@@
To enable more natural and seamless communication, recent tools’ meeting environments are designed to leverage spatial affordances that mimic social cues and collaboration strategies from face-to-face interactions. Our review surfaced five spatial composition techniques (Fig. 2): enabling interpersonal interactions through manipulating proximity and scale of users’ videos, rendering spatial audio, structuring users’ layouts to support turn-taking,
A common approach is using proxemic metaphors to facilitate conversations: in Gather444Gather: https://www.gather.town/, users can initiate one-on-one video calls by “walking up” to each other in a 2D map; OpenMic (Hu et al., 2023a) enables users to approach a “Virtual Floor” in the center of the environment to express their desire to speak. Scaling metaphors are commonly used to highlight specific users: Microsoft Teams and Zoom offer Speaker views that render active speakers at a larger size than other meeting participants; OpenMic (Hu et al., 2023a) allows users on the Virtual Floor to negotiate conversational transitions by increasing the size of their videos. To help users follow the flow of conversations, Hyrkas et al. explored requirements for rendering spatialized audio for co-located and remote users in hybrid meetings (Hyrkas et al., 2023); MirrorVerse (Grønbæk et al., 2023a) implements a “doorway” metaphor that enable users to preview conversations in breakout rooms through spatial audio. Systems like Teams TogetherMode and Ohyay place users in unified 2D environments that establish consistent user layouts (e.g., by seating users around a table), which can help establish turn-taking patterns. Perspectives (Tang et al., 2023) further supports turn-taking by rendering unique viewpoints for distributed users, which can enable more natural behaviors from co-located meetings, e.g., turning your head and making eye contact with active speakers.
@@
We observed a few examples where meeting environments can document collaborative activities, providing a basis for later iterations on ideas.
First, meeting environments can record a history of user layouts and changes to the task space, which can be replayed at a later time to understand collaboration patterns. For example, MirrorVerse (Grønbæk et al., 2023a) offers record-and-replay tools to capture changes in the entire workspace, including how users reconfigure video layouts and progressively adapt meeting controls to their needs. Finally, the meeting environment could be an artifact produced through collaborative activities, e.g., for interior design or world-building usage scenarios. In our review of prior work, we did not find examples of video-conferencing tools that explicitly claim this functionality; however, this use case is common in VR immersive authoring tools (Xia et al., 2018; Zhang et al., 2023a) which allow users to collaboratively create 3D environments while being immersed in the environment itself.
Our review of prior work surfaces a wide range of environment composition techniques that can enhance users’ awareness, communication, and sense of co-presence in distributed meetings. However, it is still a challenge for meeting participants to tailor environments to enable these affordances as needed, due to the time and manual effort required to compose environments with existing customization suites (Grønbæk et al., 2023a). With BlendScape, we contribute a flexible set of generative AI-enabled techniques that can be used to create and adapt meeting environments to serve ad-hoc collaboration activities. In Sec. 5, we return to this classification of prior video-conferencing systems to demonstrate the range of distributed collaboration techniques that BlendScape can enable.
@@ 2.2. AI-Assisted Environment Generation
Generative AI techniques have been a focal point of prior HCI research (Agarwala et al., 2004; El-Nouby et al., 2019; Dang et al., 2022; Liu et al., 2022; Dang et al., 2023), which explored the interplay between generative models and user-driven design interactions. Based on user input, such as visuals or textual prompts, these models can generate a wide variety of content including images (Avidan and Shamir, 2007; Mirza and Osindero, 2014; Zhang et al., 2017; Yu et al., 2019; Rombach et al., 2022), text (Radford et al., 2018; Brown et al., 2020; OpenAI, 2023), and even 3D objects (Coyne and Sproat, 2001; Gkioxari et al., 2020; Koh et al., 2022; Karnewar et al., 2023; Jun and Nichol, 2023).
@@
Recent advances in image generation techniques allow the generation to be conditioned in a variety of useful ways, in addition to guidance from text-based prompts. For example, ControlNet uses an auxiliary model that incorporates information conveyed through additional input modalities, such as depth, semantic, and human pose representations (Zhang et al., 2023b; Li et al., 2023). Prior work also explored intuitive techniques for end-users to control generative AI models, e.g., through sketches (Chung et al., 2022; Zhang et al., 2023b), voice (Numan et al., 2023), painting (Chung and Adar, 2023b), continuous and discrete sliders (Chung and Adar, 2023a) and iterative design mechanisms (Sharma et al., 2018; El-Nouby et al., 2019; Dang et al., 2023).
@@
There is increased interest in harnessing image generation models for the dynamic generation of virtual environments (Coyne and Sproat, 2001; Dang et al., 2023). Before the rise of AI-driven approaches, environment generation relied on predefined components and procedural logic. For instance, WordsEye converts text descriptions into corresponding 3D scenes using a fixed database of 3D models and predefined rules (Coyne and Sproat, 2001). Oasis transforms 3D-scanned physical spaces into procedurally generated virtual environments, mapping walkable areas and recognizing real-world objects to enhance immersion (Sra et al., 2018). DreamWalker used procedural techniques to substitute real-world elements with virtual ones, enabling users to stay immersed in VR while safely navigating the real world (Yang et al., 2019).
@@
These systems paved the way for the utilization of generative techniques that can be observed in recent AI-assisted creativity support tools. A recent example in this domain is WorldSmith, a system that investigates how multimodal image generation systems can be harnessed to aid users in visualizing and iteratively refining elements of fictional worlds (Dang et al., 2023). Opal (Liu et al., 2022) guides users through a structured search for visual concepts to generate images for news illustrations, utilizing GPT-3 for prompt generation including keyword, tone, icon, and style suggestions based on an article’s content.
BlendScape builds on similar techniques observed in these systems, but with a distinct focus on supporting distributed collaboration. While WorldSmith employs inpainting methods for enhancing elements in fictional 2D worlds, BlendScape utilizes inpainting to blend users’ backgrounds into unified and contextually relevant meeting environments. Like Opal, BlendScape leverages a large language model (LLM) to dynamically generate prompts aligned with user-specified meeting themes and activities.

### v2

# BlendScape: Enabling End-User Customization of Video-Conferencing Environments through Generative AI

@@ Abstract.
Today’s video-conferencing tools support a rich range of professional and social activities, but their generic meeting environments cannot be dynamically adapted to align with distributed collaborators’ needs. To enable end-user customization, we developed BlendScape, a rendering and composition system for video-conferencing participants to tailor environments to their meeting context by leveraging AI image generation techniques. BlendScape supports flexible representations of task spaces by blending users’ physical or digital backgrounds into unified environments and implements multimodal interaction techniques to steer the generation. Through an exploratory study with 15 end-users, we investigated whether and how they would find value in using generative AI to customize video-conferencing environments. Participants envisioned using a system like BlendScape to facilitate collaborative activities in the future, but required further controls to mitigate distracting or unrealistic visual elements. We implemented scenarios to demonstrate BlendScape’s expressiveness for supporting environment design strategies from prior work and propose composition techniques to improve the quality of environments.
@@ 1. Introduction
Advances in video-conferencing technologies and their increasing availability over the past few decades have enabled distributed users to collaborate on activities that previously required face-to-face interaction. Since the COVID-19 pandemic, video-conferencing has gained popularity not just for facilitating professional tasks (e.g., remote work and distance learning (Benabdallah et al., 2021; Lee et al., 2022)), but also for health appointments (Han et al., 2023), social gatherings (Fuchsberger et al., 2021), and hobbies (Cai et al., 2021; Yuan et al., 2021). However, today’s video-conferencing tools do not reflect the rich range of activities that they are used for, due to how they compose meeting environments (i.e., the “stage” or background rendered around users’ videos). Users are typically placed in separate regions of a video grid within generic meeting rooms, which can lead to meeting fatigue (Fauville et al., 2021), reduce user engagement (Choi and Diehm, 2021), and disrupt interpersonal cues for mediating conversations (Tang et al., 2023; Hu et al., 2023a).
To support more expressive video-conferencing environments that are aligned with distributed collaborators’ needs, we envision leveraging generative AI to enable end-users to create custom meeting environments. To understand the existing design space, we reviewed video-mediated communication research that redesigned meeting spaces to mitigate challenges with distributed collaboration (e.g., communication barriers, decreased sense of co-presence). We identify three main design strategies: (1) Establishing the meeting context through the environment (e.g., by rendering shared task spaces (Grønbæk et al., 2021; Hu et al., 2023b) or thematic visuals (Hunter et al., 2014; Follmer et al., 2010)); (2) Leveraging spatial metaphors to enhance communication (e.g., facilitating turn-taking via proxemic interactions between users (Hu et al., 2023a)); (3) Using the environment to record a meeting history, to aid future collaboration (Xia et al., 2018). Despite the HCI community’s knowledge of effective meeting environment designs and empirical studies demonstrating their benefits for distributed collaboration, there is a lack of tool support for end-users to implement these designs in real-time. Commercial customization tools111Microsoft Teams TogetherMode: https://www.microsoft.com/en-us/microsoft-teams/teams-together-mode; Ohyay: https://ohyay.co/ require significant manual effort, making it infeasible to adapt environments as meetings progress (Grønbæk et al., 2023a).
As a step towards this vision, we developed BlendScape, a rendering and composition system for video-conferencing participants to create environments tailored to their meeting context (Fig. 1).
@@
We introduce two key innovations: (1) We ground the generation of meeting environments in real spaces that are meaningful to users by blending their physical or virtual backgrounds into a unified environment. This can serve as a mechanism for personalization (Venolia et al., 2018) or to incorporate physical objects to collaborate around (Hu et al., 2023b; Junuzovic et al., 2012). (2) Capitalizing on recent advances in generative AI, we leverage image generation models to enable expressive and rapid techniques for composing environment designs. While such techniques are the subject of ongoing research, several dominant modes have emerged:
text-to-image: generating an image from a given text prompt.
@@
image-to-image: generating an image from a text prompt and an image prior (i.e., input image), retaining features of the image prior while introducing new elements or styles consistent with the prompt.
@@
inpainting: similar to image-to-image, but using a mask to determine which parts of the image prior should be unchanged. The rest of the image is generated in a way that it is consistent with the fixed parts of the image prior (i.e., blended).
BlendScape uses inpainting to merge users’ video backgrounds into blended environments and image-to-image techniques to transform existing images of environments to reflect the meeting purpose.
To lower the barrier for end-users to generate good quality scenes, we developed multimodal interaction techniques to steer the generation of relevant visuals and composition techniques to naturally integrate users’ videos within the scene.
We assessed the benefits and limitations of BlendScape’s customization techniques in two steps. First, to demonstrate the expressiveness, we implemented three scenarios using BlendScape, exploring a range of professional and social collaborative activities. These scenarios incorporate a majority of environment design strategies for supporting distributed collaboration from our review of prior video-conferencing systems.
@@
Second, we conducted an exploratory study with 15 end-users to investigate whether and how they would find value in using generative AI to customize video-conferencing environments. Through guiding participants to prototype meeting spaces for three scenarios using BlendScape, we elicited their customization preferences and explored to what extent BlendScape enabled them to achieve their design intentions. All participants could envision using generative AI techniques to facilitate a range of collaborative activities in the future (e.g., to spark creativity in professional settings or set a theme for social gatherings). However, to feel comfortable adapting environments during live meetings, they would require further controls to mitigate distracting or unrealistic visual elements. We propose improvements to BlendScape’s implementation to address these limitations in future work.
Our key contributions are: (1) the BlendScape composition system, which enables real-time end-user customization of video-conferencing environments through generative AI-driven composition techniques; (2) an evaluation of BlendScape’s expressiveness and considerations for empowering new design participants (Jr., 2007), through a study demonstrating how 15 video-conferencing users envision leveraging generative AI to personalize meeting environments and our implementation of three target use cases.
@@ 2. Related Work
Our work extends prior research on (1) designing video-conferencing environments to support distributed collaboration; (2) generative AI techniques for constructing 2D and 3D environments.
2.1. Design Strategies for Video-Conferencing Environments
Our goal with BlendScape was to provide a unified set of customization techniques for dynamically tailoring meeting environments to a wide range of collaborative activities. Here, we define the environment as the “stage” or background rendered around users’ videos, excluding functional tools such as the chat or host controls.
To understand the design space of meeting environments, we reviewed commercial video-conferencing tools and systems from video-mediated communication research. We classified the environment design strategies these systems adopted to mitigate challenges in distributed collaboration, as demonstrated through empirical studies. Our review surfaced three main roles that meeting environments can play in supporting distributed collaboration (Fig. 2): (1) establishing a shared context, (2) enabling spatial metaphors for communication, (3) serving as a record or artifact of collaboration.
Classification of Environment Composition Techniques. This figure shows a table with three columns. The first column has header ”Establishing a Shared Context” and children ”Rendering users’ videos within a unified environment”, ”Conveying the physical or digital task space”, ”Setting the theme or mood of a meeting. The second column as header ”Enabling Spatial Metaphors for Communication” and children ”Enabling proxemic metaphors”, ”Enabling scaling metaphors”, ”Rendering spatialized audio”, ”Establishing consistent user layouts”, ”Rendering egocentric viewing perspectives.” The third column has header ”Serving as a Record or Artifact of Collaboration” and children ”Recording a history of collaborative activities”, ”Serving as an artifact produced through collaboration.”
We focus our review on screen-based meeting systems with 2D or 2.5D designs. Design strategies for other modalities (e.g., tabletop (Zagermann et al., 2016; Tuddenham and Robinson, 2009) or mixed reality interfaces (Pejsa et al., 2016; Herskovitz et al., 2022; Grønbæk et al., 2023b)) would likely introduce new dimensions to our categorization.
@@ 2.1. Design Strategies for Video-Conferencing Environments
A primary role of meeting environments is creating a shared frame of reference for distributed users. We observed three design strategies (Fig. 2A): displaying users within a unified meeting space, incorporating elements of physical or digital task spaces, and establishing a meeting theme.
Simulating co-located meetings by rendering users’ videos within a unified environment is a popular strategy to increase distributed users’ sense of co-presence (Tang et al., 2023; Hunter et al., 2014), as demonstrated both by commercial tools (e.g., Teams Together Mode, Ohyay11footnotemark: 1) and research systems (e.g., Waazam (Hunter et al., 2014), HyperMirror (Morikawa and Maesako, 1998), BISi (O’Hara et al., 2011)). Conveying the task space, i.e., the physical or digital space where artifacts are collaboratively produced (Buxton, 2009), is critical for distributed users who would otherwise lack awareness of their collaborators’ actions. Beyond traditional screen-sharing capabilities that display digital task spaces, MirrorBlender’s (Grønbæk et al., 2021) layerable “mirrors” (translucent representations of users’ videos and shared screens) enable more natural interactions with digital artifacts, e.g., using hands to gesture around shared content. Capturing physical task spaces (e.g., for learning hands-on skills (Kumaravel et al., 2019)) often requires custom hardware setups (Junuzovic et al., 2012; Labrie et al., 2022). To mitigate this, ThingShare (Hu et al., 2023b) enables users to scan physical objects via webcams and manipulate their digital representations. To set a common theme or mood for meeting participants, Wazaam (Hunter et al., 2014) and VideoPlay (Follmer et al., 2010) introduce capture and rendering techniques for playful interactions, e.g., compositing children within storybook illustrations.
@@
To enable more natural and seamless communication, recent tools leverage spatial affordances to mimic collaboration strategies from face-to-face interactions. Our review surfaced five spatial composition techniques (Fig. 2B): enabling conversational transitions through manipulating proximity or scale of users’ videos, rendering spatial audio, structuring users’ layouts to support turn-taking, and rendering egocentric viewing perspectives.
First, meeting rooms that resemble physical spaces afford using proxemic interactions to facilitate conversations: in Gather222Gather: https://www.gather.town/, users can initiate one-on-one video calls by “walking up” to each other; in OpenMic (Hu et al., 2023a), users position themselves near a “Virtual Floor” to express their desire to speak. Scaling metaphors are commonly used to highlight specific users: Teams’ and Zoom’s Speaker views render active speakers at a larger size than other meeting participants; OpenMic (Hu et al., 2023a) allows users to negotiate conversational transitions by increasing the size of their videos. Rendering spatial audio can help users follow conversation flows and support inclusion of remote participants in hybrid meetings (Hyrkas et al., 2023; Mu et al., 2024) (e.g., MirrorVerse’s (Grønbæk et al., 2023a) “doorway” function gives users an auditory preview of breakout room discussions before they join). Maintaining consistent user layouts, (e.g., by seating users around a table in TogetherMode11footnotemark: 1) can help establish turn-taking patterns. Perspectives (Tang et al., 2023) further supports turn-taking by rendering egocentric viewpoints in the environment for each user, which simulates face-to-face social cues, e.g., making eye contact with speakers.
@@
Finally, we observed two examples where meeting environments document collaborative activities, providing a basis for later ideation (Fig. 2C). First, meeting tools can record a history of user interactions, movements, and changes to the meeting environments, which can be replayed at a later time to understand collaboration patterns (e.g., MirrorVerse’s workspace record-and-replay tools (Grønbæk et al., 2023a)). Finally, the meeting environment could be an artifact produced through users’ collaboration, e.g., for interior design or world-building scenarios. In our review of prior work, we did not find examples of video-conferencing tools that explicitly claim this functionality; however, this use case is common in VR immersive authoring tools (Xia et al., 2018; Zhang et al., 2023a) which allow users to collaboratively create virtual content while situated in the virtual world itself.
While the HCI community has significant empirical knowledge on how to configure meeting environments to enhance distributed collaboration, it is still a challenge for end-users to attain these designs, due to a lack of real-time customization support in video-conferencing tools (Grønbæk et al., 2023a).
Commercial authoring tools (e.g., from Together Mode11footnotemark: 1, Ohyay11footnotemark: 1, and Gather22footnotemark: 2) are geared towards pre-meeting use, requiring users to manually craft relevant visuals and establish user layouts. Inspired by recent customization suites like MirrorVerse (Grønbæk et al., 2023a), our work aims to lower the barrier for users to dynamically create video-conferencing environments. We leverage generative AI techniques as a new approach for rapidly expressing and aligning environment designs with meeting contexts (e.g., to visualize shared task spaces or create themed visuals).
@@ 2.2. AI-Assisted Environment Generation
Our work also builds on prior approaches for enabling non-technical users to interact with generative AI models, which has been the focus of recent HCI research (Agarwala et al., 2004; El-Nouby et al., 2019; Dang et al., 2022; Liu et al., 2022; Dang et al., 2023). Based on user input, such as visuals or text-based prompts, these models can generate a wide variety of content including images (Avidan and Shamir, 2007; Mirza and Osindero, 2014; Zhang et al., 2017; Yu et al., 2019; Rombach et al., 2022), text (Radford et al., 2018; Brown et al., 2020; OpenAI, 2023), and even 3D objects (Coyne and Sproat, 2001; Gkioxari et al., 2020; Koh et al., 2022; Karnewar et al., 2023; Jun and Nichol, 2023).
@@
Recent advances in image generation techniques allow the generation to be conditioned in a variety of useful ways, in addition to guidance from text-based prompts. For example, ControlNet uses an auxiliary model that incorporates additional data such as depth, semantic, and human pose representations (Zhang et al., 2023b; Li et al., 2023). Prior work also explored intuitive techniques for end-users to control generative AI models, e.g., through sketching (Chung et al., 2022; Zhang et al., 2023b; Chung and Adar, 2023b), speech (Numan et al., 2023), sliders (Chung and Adar, 2023a), and iterative design mechanisms (Sharma et al., 2018; El-Nouby et al., 2019; Dang et al., 2023).
@@
There is increased interest in harnessing image generation models to dynamically generate virtual environments. Before the rise of AI-driven approaches, environment generation relied on predefined components and procedural logic. For example, WordsEye converts text descriptions into 3D scenes using a database of 3D models and predefined rules (Coyne and Sproat, 2001). DreamWalker substitutes real-world elements with virtual content via procedural generation, enabling VR users to safely navigate their physical surroundings (Yang et al., 2019).
@@
These procedural approaches paved the way for the utilization of generative techniques that can be observed in recent AI-assisted creativity support tools. A recent example is WorldSmith, a system that investigates how multimodal image generation models can be harnessed to aid users in authoring and iteratively refining elements of fictional worlds (Dang et al., 2023). Opal (Liu et al., 2022) guides users through a structured search for visual concepts to generate images for news illustrations, utilizing LLMs to tune users’ prompts based on an article’s content.
BlendScape builds on similar techniques, but with a distinct focus on generating creative environments for video-conferencing systems that can be aligned with meeting participants’ goals. In particular, we employ inpainting methods to blend users’ video backgrounds into unified scenes, provide multimodal input techniques for users to steer the generation, and leverage LLMs to dynamically tailor users’ prompts to meeting themes and activities.